OpenStudy (anonymous):

−4(x + 1) − 3 = −3(x − 4)

OpenStudy (anonymous):

See, if you multiply (-4) with (x + 1) What you will get???

OpenStudy (jiteshmeghwal9):

-4(x+1)-3=-3(x-4) =>-4x-4-3=-3x+12 =>-4x-7=-3x+12 =>-4x+3x=12+7 =>-x=19 =>x=-19. U have to distribute-4 btw 'x' & '1'&-3 btw 'x' & '4'. & then solve it.

OpenStudy (jiteshmeghwal9):

CLEAR??????????